OpenStudy (anonymous):

Hey guys please help me, what are the factors of -7x+3

OpenStudy (luigi0210):

Uhm, the only thing that can really be factored is -1.. \[\LARGE -1(7x-3)\]

OpenStudy (anonymous):

hmm...actually the whole thing is 4x^2-7x+3

OpenStudy (***[isuru]***):

4x^2 - 7x + 3 = 4x^2 - 4x - 3x +3 = 4x ( x - 1) -3 ( x - 1) = ( 4x - 3 ) ( x -1) this can be factorized in the above way :)

OpenStudy (anonymous):

so what happens to 7x

OpenStudy (***[isuru]***):

if u wan to factorized it step by step.. first u need to find a pair of factor which will multiply gave 12 ( the multiplication of the constant by the coefficient of x^2) and will add up to gave 7 ( the coefficient of x)

OpenStudy (***[isuru]***):

u can c that -4 and -3 will multiply to give +12 and add up to give -7

OpenStudy (anonymous):

becuase I was looking for a pair that will give me a product of 3 and add up to give -7

OpenStudy (***[isuru]***):

nope the product should give u the multiplication of constant x ( coefficient of x^2) not just the constant (3) which is 3 x 4 = 12

OpenStudy (anonymous):

I see...thanks

OpenStudy (***[isuru]***):

u r welcome